*20 g each of NaCl(s), KCl(s), Na2NO3(s), and NH4NO3(s) are present in 500 mL of solution in four different containers.

 

Which of the following solutions has the lowest concentration of salt?


1
Expert's answer
2021-04-22T03:04:24-0400

20 g each of NaCl(s), KCl(s), Na2NO3(s), and NH4NO3(s) are present in 500 mL of solution in four different containers.

 

Which of the following solutions has the lowest concentration of salt?

Solution:

formulas that I use:

Cm=m/MV

m-mass of substance.

M-molar mass of a substance

V-solvent volume

А) for NaCl

M=Ar(Na)+Ar(Cl)=23+35.5=58.5g/mol

Solvent volume=500mL=0.5L

Cm=20/58.5x0.5=0.6838 mol/l

b)For KCl

M(KCL)=39+35.5=74.5 g/mol

Cm=20/74.5x0.5=0.5369 mol/L

c) For Na2NO3

M(Na2NO3)=23x2+14+16x3=108 g/mol

Cm=20/108x0.5=0.37mol/L

d)For NH4NO3

M(NH4NO3)=14+1x4+14+16x3=80 g/mol

Cm=20/80x0.5=0.5 mol/L

Answer:

0.37<0.5<0.5369<0.6838

Na2NO3 has the lowest concentration of these salts